OpenStudy (anonymous):

Find the sum: (3x^2+5x-8) + (5x^2-13x-5) Help plspls

OpenStudy (anonymous):

For sum you can sum same ones ! same have same words and powers . for example : X^2 and 2X^2 are same . x and x^2 are not . in this question you can sum : 3x^2 & 5x^2 +5x & -13x -8 & -5 Together ;) (3x^2+5x-8) + (5x^2-13x-5) = (3x^2+5x^2)+(5x-13x)+(-5-8) = (3+5)x^2+(5-13)x-5-8 can you answer now ?!;)

OpenStudy (anonymous):

 Ohlorrd. tr||jahsjdkfhkjsad. Thankurr.

OpenStudy (anonymous):

ok ! (3+5)x^2+(5-13)x-5-8 =8x^2-8x-13